# Reinforced In-Context Black-Box Optimization

### v1

@@ Abstract
Black-Box Optimization (BBO) has found successful applications in many fields of science and engineering. Recently, there has been a growing interest in meta-learning particular components of BBO algorithms to speed up optimization and get rid of tedious hand-crafted heuristics. As an extension, learning the entire algorithm from data requires the least labor from experts and can provide the most flexibility. In this paper, we propose RIBBO, a method to reinforce-learn a BBO algorithm from offline data in an end-to-end fashion. RIBBO employs expressive sequence models to learn the optimization histories produced by multiple behavior algorithms and tasks, leveraging the in-context learning ability of large models to extract task information and make decisions accordingly. Central to our method is to augment the optimization histories with regret-to-go tokens, which are designed to represent the performance of an algorithm based on cumulative regret of the histories. The integration of regret-to-go tokens enables RIBBO to automatically generate sequences of query points that satisfy the user-desired regret, which is verified by its universally good empirical performance on diverse problems, including BBOB functions, hyper-parameter optimization and robot control problems.
Machine Learning, ICML
@@ 1 Introduction
Black-Box Optimization (BBO) [4; 1] refers to optimizing objective functions where neither analytic expressions nor derivatives of the objective are available. To solve BBO problems, we can only access the results of objective evaluation, which usually also incurs a high computational cost. Many fundamental problems in science and engineering involve optimization of expensive BBO functions, such as drug discovery [41; 52], material design [21; 23], robot control [10; 11], and optimal experimental design [24; 43], just to name a few.
@@
To date, a lot of BBO algorithms have been developed, among which the most prominent ones are Bayesian Optimization (BO) [50; 20] and Evolutionary Algorithms (EA) [5; 65]. Despite the advancements, these algorithms typically solve BBO problems from scratch and rely on expert-derived heuristics. Consequently, they are often hindered by slow convergence rates, and unable to leverage the inherent structures within the optimization problems [3; 6].
@@
Recently, there has been a growing interest in meta-learning a particular component of the algorithms with previously collected data [18; 2]. Learning the component not only alleviates the need for the laborious design process of the domain experts, but also specifies the component with domain data to facilitate subsequent optimization. For example, some components in BO are proposed to be learned from data, including the surrogate model [45; 58; 61; 40], acquisition function [56; 29], initialization strategy [17; 46], and search space [44; 36]; some core evolutionary operations in EA have also been considered, e.g., learning the selection and mutation rate adaptation in genetic algorithm [33] or the update rules for evolution strategy [34].
@@
There have also been some attempts to learn an entire algorithm in an End-to-End (E2E) fashion, which requires almost no expert knowledge at all and provides the most flexibility across a broad range of BBO problems. However, existing practices require additional knowledge regarding the objective function during the training stage, e.g., the gradient information (often impractical for BBO) [13; 53] or online sampling from the objective function (often very expensive) [38]Chen et al. [14] proposed the OptFormer method to imitate the behavior algorithms separately during training, requiring a user to manually specify which algorithm to execute during testing, which still poses a challenge. Thus, these methods are less ideal for practical scenarios where offline datasets are often available beforehand and a suitable algorithm for the given task has to be identified automatically without the involvement of domain experts.
@@
In this paper, we introduce Reinforced In-context BBO (RIBBO), which learns a reinforced BBO algorithm from offline datasets in an E2E fashion. RIBBO employs an expressive sequence model, i.e., causal transformer, to fit the optimization histories in the offline datasets generated by executing multiple behavior algorithms on multiple tasks. The sequence model is fed with previous query points and their function values, and trained to predict the distribution over the next query point. During testing, the sequence model itself serves as a BBO algorithm by generating the query points auto-regressively. Apart from this, RIBBO augments the optimization histories with regret-to-go (RTG) tokens, which are calculated by summing up the regrets over the future part of the histories, representing the future performance of an algorithm. By integrating the RTG tokens into the modeling, RIBBO can automatically identify different algorithms, and generate sequences of query points that satisfy the specified regret. Such modeling enables RIBBO to circumvent the impact of inferior data and further reinforce its performance on top of the behavior algorithms.
@@ 2.1 Black-Box Optimization
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a black-box function, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a $d$-dimensional search space. The goal of BBO is to find an optimal solution $\bm{x}^{*}\in\mathop{\arg\max}_{\bm{x}\in\mathcal{X}}f(\bm{x})$, with the only permission of querying the objective function value. Several classes of BBO algorithms have been proposed, e.g., BO [50; 20] and EA [5; 65]. The basic framework of BO contains two critical components: a surrogate model, typically formalized as Gaussian Process (GP) [48], and an acquisition function [60], which are used to model $f$ and decide the next query point, respectively. EA is a class of heuristic optimization algorithms inspired by natural evolution. It maintains a population of solutions and iterates through mutation, crossover, and selection operations to find better solutions.
@@ 2.2 Meta-Learning in Black-Box Optimization
Hand-crafted BBO algorithms usually require an expert to analyze the algorithms’ behavior on a wide range of problems, which is tedious and time-consuming. One solution is meta-learning [55; 28], which aims to exploit knowledge to improve the performance of learning algorithms given data from a set of tasks. By parameterizing a component of BBO algorithms or even an entire algorithm that is traditionally manually designed, we can learn from historical data to incorporate the knowledge into the optimization, which may bring speedup.
@@
Meta-learning particular components has been studied with different BBO algorithms. Meta-learning in BO can be divided into four main categories according to “what to transfer” [6], including the design of the surrogate model, acquisition function, initialization strategy, and search space. For surrogate model design, Wang et al. [58]; Wistuba & Grabocka [61] parameterized the mean or kernel function of the GP model with Multi-Layer Perceptron (MLP), while Perrone et al. [45]; Müller et al. [40] substituted GP with Bayesian linear regression or neural process [22; 39]. For acquisition function design, MetaBO [56] uses Reinforcement Learning (RL) to meta-train an acquisition function on a set of related tasks, and FSAF [29] employs a Bayesian variant of deep Q-network as a surrogate differentiable acquisition function trained by model-agnostic meta-learning [19]. The remaining two categories focus on exploiting the previous good solutions to warm start the optimization [17; 46] or shrink the search space [44; 36]. Meta-learning in EA usually focuses on specific evolutionary operations. For example, Lange et al. [33] substituted core genetic operators, including selection and mutation rate adaptation, with dot-product attention modules; Lange et al. [34] meta-learned a self-attention-based architecture to discover effective and order-invariant update rules.
@@
Meta-learning entire algorithms has also been explored to obtain more flexible models. Early works [13; 53] uses Recurrent Neural Network (RNN) to meta-learn a BBO algorithm by optimizing the summed objective functions of some iterations. RNN uses its memory state to store information about history and outputs the next query point. This work assumes access to gradient information during the training phase, which is, however, usually impractical in BBO problems. OptFormer [14] uses a text-based transformer framework to learn an algorithm, providing a universal E2E interface for BBO problems. It can simultaneously imitate at least $7$ different BBO algorithms on a broad range of problems. However, when using OptFormer to solve problems, a user must specify an algorithm manually. How to select suitable algorithms automatically is still a challenge. Neural Acquisition Processes (NAP) [38] uses transformer to meta-learn the surrogate model and acquisition function of BO jointly. Due to the lack of labeled acquisition data, NAP uses an online RL algorithm with a supervised auxiliary loss for training, which requires online sampling from the expensive objective function and lacks efficiency. Compared to these two state-of-the-art E2E methods, i.e., OptFormer [14] and NAP [38], our approach offers the advantage of automatically identifying and deploying the best-performing algorithm without requiring the user to pre-specify which algorithm to use during the test phase, and utilizes a supervised learning loss for training on a fixed offline dataset without the need for further interaction with the objective function.
@@ 2.3 Decision Transformer
Transformer has emerged as a powerful architecture for sequence modeling tasks [62; 32; 59]. A basic building block behind transformer is the self-attention mechanism [54], which captures the correlation between tokens of any pair of timesteps. As the scale of data and model increases, transformer has demonstrated the in-context learning ability [9], which refers to the capability of the model to infer the tasks at hand based on the input contexts.
Decision Transformer (DT) [12] abstracts RL as a sequence modeling problem, and introduces return-to-go tokens, representing the cumulative rewards over future interactions. Conditioning on return-to-go tokens enables DT to correlate the trajectories with their corresponding returns and generate future actions to achieve a user-specified return. Inspired by DT, we will treat BBO tasks as a sequence modeling problem naturally, use a causal transformer for modeling, and train it by conditioning on future regrets. Such design is expected to enable the learned model to distinguish algorithms with different performance and achieve good performance with a user-specified low regret.
@@
Figure 1: Illustration of RIBBO. Left: Data Generation. $K$ existing BBO algorithms $\{\mathcal{A}_{j}\}_{j=1}^{K}$ and $N$ BBO tasks $\{f_{i}\}_{i=1}^{N}$ are used to serve as the behavior algorithms and the training tasks, respectively. The offline datasets $\{\mathcal{D}_{i,j}\}$ consist of the optimization histories $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ collected by executing each behavior algorithm $\mathcal{A}_{j}$ on each task $f_{i}$ for $T$ evaluation steps, which are then augmented with the regret-to-go tokens $R_{t}$ (calculated as the cumulative regret $\sum^{T}_{t^{\prime}=t+1}(y^{*}-y_{t^{\prime}})$ over the future optimization history) to generate the final dataset $\{\widehat{\mathcal{D}}_{i,j}\}$ for training. Right: Training and Inference. Our model takes in triplets of $(\bm{x}_{t},y_{t},R_{t})$, embeds them into one token, and outputs the distribution over the next query point $\bm{x}_{t+1}$. During training, the ground-truth next query is used to minimize the loss in Eq. (4). During inference, the next query point $\bm{x}_{t+1}$ is generated auto-regressively based on the current history $\hat{\bm{h}}_{t}$.
@@ 3 Method
This section presents Reinforced In-context Black-Box Optimization (RIBBO), which learns an enhanced BBO algorithm in an E2E fashion. The overall workflow is presented in Figure 1.
3.1 Problem Formulation
We follow the task-distribution assumption, which is commonly adopted in meta-learning settings [19; 28; 64]. Our goal is to learn a generalizable model $\mathcal{M}$ capable of solving a wide range of BBO tasks, each associated with a BBO objective function $f$ sampled from the task distribution $P(\mathcal{F})$, where $\mathcal{F}$ denotes the function space.
Let $[N]$ denote the integer set $\{1,2,\ldots,N\}$. During training, we usually access $N$ source tasks and each task corresponds to an objective function $f_{i}\sim P(\mathcal{F})$, where $i\in[N]$. Hereafter, we use $f_{i}$ to denote the task $i$ if the context is clear. We assume that the information is available via offline datasets $\mathcal{D}_{i,j}$, which are produced by executing a behavior algorithm $\mathcal{A}_{j}$ on task $f_{i}$, where $j\in[K]$ and $i\in[N]$. Each dataset $\mathcal{D}_{i,j}=\{\bm{h}^{i,j,m}_{T}\}_{m=1}^{M}$ consists of $M$ optimization histories $\bm{h}^{i,j,m}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$, where $\bm{x}_{t}$ is the query point selected by $\mathcal{A}_{j}$ at iteration $t$, and $y_{t}=f_{i}(\bm{x}_{t})$ is its objective value. If the context is clear, we will omit $i,j,m$ and simply use $\bm{h}_{T}$ to denote an optimization history with length $T$. The initial history $\bm{h}_{0}$ is defined as $\emptyset$. We impose no additional assumptions about the behavior algorithms, allowing for a range of BBO algorithms, even random search.
@@ 3.1 Problem Formulation
as the evaluation metric, where $y^{*}$ is the optimum value and $\{y_{t}\}^{T}_{t=1}$ are the function values of the points in $\bm{h}_{T}$.
@@ 3.2 Method Outline
Given the current history $\bm{h}_{t-1}$ at iteration $t$, a BBO algorithm selects the next query point $\bm{x}_{t}$, observes the function value $y_{t}=f_{i}(\bm{x}_{t})$, and updates the history $\bm{h}_{t}=\bm{h}_{t-1}\cup\{(\bm{x}_{t},y_{t})\}$. Similar to the previous work [13], we take this framework as a starting point and define the learning of a universal BBO algorithm as learning a model $\mathcal{M}_{\bm{\theta}}$, which takes the preceding history $\bm{h}_{t-1}$ as input and outputs a distribution of the next query point $\bm{x}_{t}$. The optimization histories in offline datasets provide a natural supervision for learning.
Suppose we have a set of histories $\{\bm{h}_{T}\}$, generated by a single behavior algorithm $\mathcal{A}$ on a single task $f$. By employing a causal transformer model $\mathcal{M}_{\bm{\theta}}$, we expect $\mathcal{M}_{\bm{\theta}}$ to imitate $\mathcal{A}$ and produce similar optimization history on $f$. In practice, we usually have datasets containing histories from multiple behavior algorithms $\{\mathcal{A}_{j}\}_{j=1}^{K}$ on multiple tasks $\{f_{i}\}_{i=1}^{N}$. To fit $\mathcal{M}_{\bm{\theta}}$, we use the negative log-likelihood loss
$\mathcal{L}_{\rm BC}(\bm{\theta})=-\mathbb{E}_{\bm{h}_{T}\sim\mathcal{D}_{i,j}}\left[\sum_{t=1}^{T}\log\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\bm{h}_{t-1})\right].$ (2)
@@
$\mathcal{L}_{\rm{RIBBO}}(\bm{\theta})=-\mathbb{E}_{\hat{\bm{h}}_{T}\sim\widehat{\mathcal{D}}_{i,j}}\left[\sum_{t=1}^{T}\log\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\hat{\bm{h}}_{t-1})\right].$ (4)
The integration of RTG tokens offers several advantages. [label=0)] 1. RTG tokens in the context bring identifiability of behavior algorithms, and the model $\mathcal{M}_{\bm{\theta}}$ can effectively utilize them to make appropriate decisions. 2. RTG tokens have a direct correlation with the metric of interest, i.e., cumulative regret $\mathrm{Reg}_{T}$ in Eq. (1). Conditioning on a lower RTG token provides a guidance to our model and reinforces $\mathcal{M}_{\bm{\theta}}$ to exhibit superior performance. These advantages will be clearly shown by experiments in Section 4.4.
Note that the resulting method RIBBO has implicitly utilized the in-context learning capacity of transformer to guide the optimization with previous histories and the desired future regret as context. The in-context learning capacity of inferring the tasks at hand based on the input contexts has been observed as the scale of data and model increases [31]. It has been explored to infer general functional relationships as supervised learning or RL algorithms. For example, Guo et al. [25]; Hollmann et al. [27]; Li et al. [37] feed the training dataset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{t-1}$ as the context and expect the model to make an accurate prediction on the query input $\bm{x}_{t}$askin et al. [35] learn RL algorithms using causal transformers. Here, we use it for BBO.

### v2

@@ Abstract
Black-Box Optimization (BBO) has found successful applications in many fields of science and engineering. Recently, there has been a growing interest in meta-learning particular components of BBO algorithms to speed up optimization and get rid of tedious hand-crafted heuristics. As an extension, learning the entire algorithm from data requires the least labor from experts and can provide the most flexibility. In this paper, we propose RIBBO, a method to reinforce-learn a BBO algorithm from offline data in an end-to-end fashion. RIBBO employs expressive sequence models to learn the optimization histories produced by multiple behavior algorithms and tasks, leveraging the in-context learning ability of large models to extract task information and make decisions accordingly. Central to our method is to augment the optimization histories with regret-to-go tokens, which are designed to represent the performance of an algorithm based on cumulative regret over the future part of the histories. The integration of regret-to-go tokens enables RIBBO to automatically generate sequences of query points that satisfy the user-desired regret, which is verified by its universally good empirical performance on diverse problems, including BBO benchmark functions, hyper-parameter optimization and robot control problems.
@@ 1 Introduction
Black-Box Optimization (BBO) [1, 4] refers to optimizing objective functions where neither analytic expressions nor derivatives of the objective are available. To solve BBO problems, we can only access the results of objective evaluation, which usually also incurs a high computational cost. Many fundamental problems in science and engineering involve optimization of expensive BBO functions, such as drug discovery [42, 55], material design [21, 23], robot control [10, 11], and optimal experimental design [24, 44], just to name a few.
@@
To date, a lot of BBO algorithms have been developed, among which the most prominent ones are Bayesian Optimization (BO) [20, 51] and Evolutionary Algorithms (EA) [5, 68]. Despite the advancements, these algorithms typically solve BBO problems from scratch and rely on expert-derived heuristics. Consequently, they are often hindered by slow convergence rates, and unable to leverage the inherent structures within the optimization problems [3, 6].
@@
Recently, there has been a growing interest in meta-learning a particular component of the algorithms with previously collected data [2, 18]. Learning the component not only alleviates the need for the laborious design process of the domain experts, but also specifies the component with domain data to facilitate subsequent optimization. For example, some components in BO are proposed to be learned from data, including the surrogate model [41, 46, 61, 64], acquisition function [29, 59], initialization strategy [17, 47], and search space [37, 45]; some core evolutionary operations in EA have also been considered, e.g., learning the selection and mutation rate adaptation in genetic algorithm [34] or the update rules for evolution strategy [35].
@@
There have also been some attempts to learn an entire algorithm in an End-to-End (E2E) fashion, which requires almost no expert knowledge at all and provides the most flexibility across a broad range of BBO problems. However, existing practices require additional knowledge regarding the objective function during the training stage, e.g., the gradient information (often impractical for BBO) [13, 56] or online sampling from the objective function (often very expensive) [39]Chen et al. [14] proposed the OptFormer method to imitate the behavior algorithms separately during training, requiring a user to manually specify which algorithm to execute during testing, which still poses a challenge. Thus, these methods are less ideal for practical scenarios where offline datasets are often available beforehand and a suitable algorithm for the given task has to be identified automatically without the involvement of domain experts.
@@
In this paper, we introduce Reinforced In-context BBO (RIBBO), which learns a reinforced BBO algorithm from offline datasets in an E2E fashion. RIBBO employs an expressive sequence model, i.e., causal transformer, to fit the optimization histories in the offline datasets generated by executing diverse behavior algorithms on multiple tasks. The sequence model is fed with previous query points and their function values, and trained to predict the distribution over the next query point. During testing, the sequence model itself serves as a BBO algorithm by generating the query points auto-regressively. Apart from this, RIBBO augments the optimization histories with regret-to-go (RTG) tokens, which are calculated by summing up the regrets over the future part of the histories, representing the future performance of an algorithm. A novel Hindsight Regret Relabelling (HRR) strategy is proposed to update the RTG tokens during testing. By integrating the RTG tokens into the modeling, RIBBO can automatically identify different algorithms, and generate sequences of query points that satisfy the specified regret. Such modeling enables RIBBO to circumvent the impact of inferior data and further reinforce its performance on top of the behavior algorithms.
@@ 2.1 Black-Box Optimization
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a black-box function, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a $d$-dimensional search space. The goal of BBO is to find an optimal solution $\bm{x}^{*}\in\mathop{\arg\max}_{\bm{x}\in\mathcal{X}}f(\bm{x})$, with the only permission of querying the objective function value. Several classes of BBO algorithms have been proposed, e.g., BO [20, 51] and EA [5, 68]. The basic framework of BO contains two critical components: a surrogate model, typically formalized as Gaussian Process (GP) [49], and an acquisition function [63], which are used to model $f$ and decide the next query point, respectively. EA is a class of heuristic optimization algorithms inspired by natural evolution. It maintains a population of solutions and iterates through mutation, crossover, and selection operations to find better solutions.
@@ 2.2 Meta-Learning in Black-Box Optimization
Hand-crafted BBO algorithms usually require an expert to analyze the algorithms’ behavior on a wide range of problems, which is tedious and time-consuming. One solution is meta-learning [28, 58], which aims to exploit knowledge to improve the performance of learning algorithms given data from a set of tasks. By parameterizing a component of BBO algorithms or even an entire algorithm that is traditionally manually designed, we can learn from historical data to incorporate the knowledge into the optimization, which may bring speedup.
@@
Meta-learning particular components has been studied with different BBO algorithms. Meta-learning in BO can be divided into four main categories according to “what to transfer” [6], including the design of the surrogate model, acquisition function, initialization strategy, and search space. For surrogate model design, Wang et al. [61] and Wistuba and Grabocka [64] parameterized the mean or kernel function of the GP model with Multi-Layer Perceptron (MLP), while Perrone et al. [46] and Müller et al. [41] substituted GP with Bayesian linear regression or neural process [22, 40]. For acquisition function design, MetaBO [59] uses Reinforcement Learning (RL) to meta-train an acquisition function on a set of related tasks, and FSAF [29] employs a Bayesian variant of deep Q-network as a surrogate differentiable acquisition function trained by model-agnostic meta-learning [19]. The remaining two categories focus on exploiting the previous good solutions to warm start the optimization [17, 47] or shrink the search space [37, 45]. Meta-learning in EA usually focuses on specific evolutionary operations. For example, Lang et al. substituted core genetic operators, i.e., selection and mutation rate adaptation, with dot-product attention modules [34], and meta-learned a self-attention-based architecture to discover effective and order-invariant update rules [35].
@@
Meta-learning entire algorithms has also been explored to obtain more flexible models. Early works [13, 56] use Recurrent Neural Network (RNN) to meta-learn a BBO algorithm by optimizing the summed objective functions of some iterations. RNN uses its memory state to store information about history and outputs the next query point. This work assumes access to gradient information during the training phase, which is, however, usually impractical in BBO. OptFormer [14] uses a text-based transformer framework to learn an algorithm, providing a universal E2E interface for BBO problems. It can simultaneously imitate at least $7$ different BBO algorithms on a broad range of problems, which, however, requires a user to specify an algorithm manually when applying it to solve problems. Thus, how to select suitable algorithms automatically is still a challenge. Neural Acquisition Processes (NAP) [39] uses transformer to meta-learn the surrogate model and acquisition function of BO jointly. Due to the lack of labeled acquisition data, NAP uses an online RL algorithm with a supervised auxiliary loss for training, which requires online sampling from the expensive objective function and lacks efficiency. Black-box Optimization NETworks (BONET) [33] employ a transformer model to fit regret-augmented trajectories in an offline BBO scenario, where the training and testing data are from the same objective function, and a prefix sequence is required to warm up the optimization before testing. Compared to the above state-of-the-art E2E methods, we consider the meta-BBO setting, where the training datasets consist of diverse algorithms across different functions. Our approach offers the advantage of automatically identifying and deploying the best-performing algorithm without requiring the user to pre-specify which algorithm to use or to provide a prefix sequence during the testing phase, and utilizes a supervised learning loss for training on a fixed offline dataset without the need for further interaction with the objective function.
@@ 2.3 Decision Transformer
Transformer has emerged as a powerful architecture for sequence modeling tasks [32, 62, 65]. A basic building block behind transformer is the self-attention mechanism [57], which captures the correlation between tokens of any pair of timesteps. As the scale of data and model increases, transformer has demonstrated the in-context learning ability [9], which refers to the capability of the model to infer the tasks at hand based on the input contexts. Decision Transformer (DT) [12] abstracts RL as a sequence modeling problem, and introduces return-to-go tokens, representing the cumulative rewards over future interactions. Conditioning on return-to-go tokens enables DT to correlate the trajectories with their corresponding returns and generate future actions to achieve a user-specified return. Inspired by DT, we will treat BBO tasks as a sequence modeling problem naturally, use a causal transformer for modeling, and train it by conditioning on future regrets. Such design is expected to enable the learned model to distinguish algorithms with different performance and achieve good performance with a user-specified low regret.
@@
Figure 1: Illustration of RIBBO. Left: Data Generation. $K$ existing BBO algorithms $\{\mathcal{A}_{j}\}_{j=1}^{K}$ and $N$ BBO tasks $\{f_{i}\}_{i=1}^{N}$ are used to serve as the behavior algorithms and the training tasks, respectively. The offline datasets $\{\mathcal{D}_{i,j}\}$ consist of the optimization histories $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ collected by executing each behavior algorithm $\mathcal{A}_{j}$ on each task $f_{i}$ for $T$ evaluation steps, which are then augmented with the regret-to-go tokens $R_{t}$ (calculated as the cumulative regret $\sum^{T}_{t^{\prime}=t+1}(y^{*}-y_{t^{\prime}})$ over the future optimization history) to generate the final dataset $\{\widehat{\mathcal{D}}_{i,j}\}$ for training. Right: Training and Inference. Our model takes in triplets of $(\bm{x}_{t},y_{t},R_{t})$, embeds them into one token, and outputs the distribution over the next query point $\bm{x}_{t+1}$. During training, the ground-truth next query point is used to minimize the loss in Eq. (4). During inference, the next query point $\bm{x}_{t+1}$ is generated auto-regressively based on the current history $\hat{\bm{h}}_{t}$.
@@ 3 Method
This section presents Reinforced In-context Black-Box Optimization (RIBBO), which learns an enhanced BBO algorithm in an E2E fashion, as illustrated in Figure 1. We follow the task-distribution assumption, which is commonly adopted in meta-learning settings [19, 28, 67]. Our goal is to learn a generalizable model $\mathcal{M}$ capable of solving a wide range of BBO tasks, each associated with a BBO objective function $f$ sampled from the task distribution $P(\mathcal{F})$, where $\mathcal{F}$ denotes the function space.
Let $[N]$ denote the integer set $\{1,2,\ldots,N\}$. During training, we usually access $N$ source tasks and each task corresponds to an objective function $f_{i}\sim P(\mathcal{F})$, where $i\in[N]$. Hereafter, we use $f_{i}$ to denote the task $i$ if the context is clear. We assume that the information is available via offline datasets $\mathcal{D}_{i,j}$, which are produced by executing a behavior algorithm $\mathcal{A}_{j}$ on task $f_{i}$, where $j\in[K]$ and $i\in[N]$. Each dataset $\mathcal{D}_{i,j}=\{\bm{h}^{i,j,m}_{T}\}_{m=1}^{M}$ consists of $M$ optimization histories $\bm{h}^{i,j,m}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$, where $\bm{x}_{t}$ is the query point selected by $\mathcal{A}_{j}$ at iteration $t$, and $y_{t}=f_{i}(\bm{x}_{t})$ is its objective value. If the context is clear, we will omit $i,j,m$ and simply use $\bm{h}_{T}$ to denote a history with length $T$. The initial history $\bm{h}_{0}$ is defined as $\emptyset$. We impose no additional assumptions about the behavior algorithms, allowing for a range of BBO algorithms, even random search.
@@
as the evaluation metric, where $y^{*}$ is the optimum value and $\{y_{t}\}^{T}_{t=1}$ are the function values in $\bm{h}_{T}$.
@@ 3.1 Method Outline
Given the current history $\bm{h}_{t-1}$ at iteration $t$, a BBO algorithm selects the next query point $\bm{x}_{t}$, observes the function value $y_{t}=f_{i}(\bm{x}_{t})$, and updates the history $\bm{h}_{t}=\bm{h}_{t-1}\cup\{(\bm{x}_{t},y_{t})\}$. Similar to the previous work [13], we take this framework as a starting point and treat the learning of a universal BBO algorithm as learning a model $\mathcal{M}_{\bm{\theta}}$, which takes the preceding history $\bm{h}_{t-1}$ as input and outputs a distribution of the next query point $\bm{x}_{t}$. The optimization histories in offline datasets provide a natural supervision for learning.
Given a set of histories $\{\bm{h}_{T}\}$ generated by a single behavior algorithm $\mathcal{A}$ on a single task $f$, we employ a causal transformer model $\mathcal{M}_{\bm{\theta}}$, and expect it to imitate $\mathcal{A}$ and produce similar optimization history on $f$. In practice, we usually have datasets containing histories from multiple behavior algorithms $\{\mathcal{A}_{j}\}_{j=1}^{K}$ on multiple tasks $\{f_{i}\}_{i=1}^{N}$. To fit $\mathcal{M}_{\bm{\theta}}$, we use the negative log-likelihood loss
$\mathcal{L}_{\rm BC}(\bm{\theta})=-\mathbb{E}_{\bm{h}_{T}\sim\mathcal{D}_{i,j}}\left[\sum\nolimits_{t=1}^{T}\log\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\bm{h}_{t-1})\right].$ (2)
@@
$\mathcal{L}_{\rm{RIBBO}}(\bm{\theta})=-\mathbb{E}_{\hat{\bm{h}}_{T}\sim\widehat{\mathcal{D}}_{i,j}}\left[\sum\nolimits_{t=1}^{T}\log\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\hat{\bm{h}}_{t-1})\right].$ (4)
The integration of RTG tokens in the context brings identifiability of behavior algorithms, and the model $\mathcal{M}_{\bm{\theta}}$ can effectively utilize them to make appropriate decisions. Furthermore, RTG tokens have a direct correlation with the metric of interest, i.e., cumulative regret $\mathrm{Reg}_{T}$ in Eq. (1). Conditioning on a lower RTG token provides a guidance to our model and reinforces $\mathcal{M}_{\bm{\theta}}$ to exhibit superior performance. These advantages will be clearly shown by experiments in Section 4.4.
The resulting method RIBBO has implicitly utilized the in-context learning capacity of transformer to guide the optimization with previous histories and the desired future regret as context. The in-context learning capacity of inferring the tasks at hand based on the input contexts has been observed as the scale of data and model increases [31]. It has been explored to infer general functional relationships as supervised learning or RL algorithms. For example, the model is expected to predict accurately on the query input $\bm{x}_{t}$ by feeding the training dataset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{t-1}$ as the context [25, 27, 38]Laskin et al. [36] learned RL algorithms using causal transformers. Here, we use it for BBO.